# Emergent Specialization: Rare Token Neurons in Language Models

### v1

@@ 1 Introduction
Recent advances in mechanistic interpretability have developed various techniques for analyzing individual neurons within transformer models. While studies have revealed that neurons encode interpretable features ranging from syntactic relationships [22, 12] to semantic concepts [14, 6], most work has focused on common patterns rather than specialized mechanisms for rare events. Notable exceptions include work Stolfo et al. [33] discovering neurons that modulate token logits proportionally to frequency. We extend their investigate whether individual neurons in the final MLP layer of transformer-based language models spontaneously specialize for rare token processing during training.
@@ 3.2 Co-activation Patterns Through the Lens of Activation Space Geometry
Construction of the activation space
Two geometric statistics
3.2.1 Activation Correlation
$D_{ij}=1-|\rho_{ij}|,$ (4)
3.3 Distribution of Neuron Influence and Phase Transitions
@@ 5 Discussions and Conjectures
Hypothesis 5.1 (Dual-Regime Organization)
@@
Hypothesis 5.2 (Parallel Mechanism Hypothesis)
@@
The parallel mechanism hypothesis suggests that rare token processing in language models involves both distributed computation (power-law phase) and specialized neuron group (plateau phase). This dual-system architecture resembles the complementary learning systems (CLS) observed in the human memory system [26, 18], where general statistical learning occurs alongside specialized mechanisms for handling exceptional cases. Analogously, the plateau neurons might function as a specialized memory system for encoding rare linguistic patterns that would otherwise be overwhelmed by the statistics of other tokens.
@@ 6 Conclusion
Based on these findings, we proposed the Dual-Regime Organization hypothesis, suggesting qualitatively different computational regimes across neuron groups, and the Parallel Mechanism hypothesis, positing that rare token processing involves both distributed and specialized computation analogous to complementary learning systems in biological memory.

### v2

@@ 1 Introduction
Recent advances in mechanistic interpretability have developed various techniques for analyzing individual neurons within transformer models. While studies have revealed that neurons encode interpretable features ranging from syntactic relationships [22, 12] to semantic concepts [14, 6], most work has focused on common patterns rather than specialized mechanisms for rare events. Notable exceptions include work Stolfo et al. [33] discovering neurons that modulate token logits proportionally to frequency. In this study, we focus on decoder-only Transformer-based models and extend their work to focus on rare tokens and investigate how individual neurons in the final MLP layer of transformer-based language models spontaneously specialize in processing these tokens during training.
@@ 3.1 Rare Token Neuron Identification
3.2 Distribution and Structure of Neuron Influence
@@ 3.3 Co-activation Patterns Through the Lens of Activation Space Geometry
Construction of the activation space
Two geometric statistics for co-activation detection
Clustering and activation correlations
$D_{ij}=1-|\rho_{ij}|,$ (6)
@@ 5 Discussions and Conjectures
Conjecture 5.1 (Dual-Regime Organization)
@@
Conjecture 5.2 (Parallel Mechanism Conjecture)
@@
The parallel mechanism conjecture suggests that rare token processing in language models involves both distributed computation (power-law phase) and specialized neuron group (plateau phase). This dual-system architecture resembles the complementary learning systems (CLS) observed in the human memory system [26, 18], where general statistical learning occurs alongside specialized mechanisms for handling exceptional cases. Analogously, the plateau neurons might function as a specialized memory system for encoding rare linguistic patterns that would otherwise be overwhelmed by the statistics of other tokens.
@@ 6 Conclusion
Based on these findings, we proposed the Dual-Regime Organization conjecture, suggesting qualitatively different computational regimes across neuron groups, and the Parallel Mechanism conjecture, positing that rare token processing involves both distributed and specialized computation analogous to complementary learning systems in biological memory.
@@
Appendix A Appendix
A.1 Limitations
Our study presents several important limitations that warrant acknowledgment. First, our analysis focuses exclusively on neurons in the final MLP layer, while rare token processing likely involves distributed mechanisms across multiple architectural components throughout the model. Future work examining attention heads, intermediate layers, and cross-layer interactions could provide a more comprehensive understanding of rare token processing mechanisms.
Second, our methodology relies on ablation-based proxies, particularly change in loss, to quantify individual neuron contributions. While these measures provide useful insights, they lack the theoretical rigor of more principled attribution methods. Developing precise theoretical frameworks for measuring neuron-level effects would strengthen our mechanistic interpretations and enable more robust conclusions about functional specialization.
Finally, our analysis is constrained to next-token prediction in autoregressive language modeling contexts. The generalizability of our findings to downstream applications—including question-answering, mathematical reasoning, and domain-specific tasks—remains an open question. Investigating rare token processing mechanisms in these applied settings would illuminate their practical significance for real-world model performance.
A.2 Results
Activation Correlation Analysis
We examined activation patterns across neuron groups in five pre-trained language models, comparing rare token neurons (boost and suppress groups) against random baseline controls. Table 1 presents pairwise activation correlations within and between neuron groups (group size = 50).
The results demonstrate consistently higher intra-group correlations for both boost neurons (range: 0.004–0.036) and suppress neurons (range: 0.011–0.052) compared to random controls (range: -0.007–0.017). Notably, cross-group correlations between boost and suppress neurons (B vs. S) show positive values (0.010–0.040), suggesting coordinated but potentially antagonistic functionality. Random baseline comparisons (R1 vs. R2) exhibit near-zero correlations, confirming the specificity of our identified neuron groups.
Table 1: Pairwise activation correlations within and between neuron groups
Model Size Boost Suppress Random B vs. R S vs. R B vs. S R1 vs. R2
Pythia-70M 70M 0.028 0.045 0.011 0.002 0.005 0.027 0.009
Pythia-410M 410M 0.036 0.052 0.007 0.005 0.006 0.040 0.007
GPT2-Small 124M 0.017 0.019 0.017 -0.001 -0.001 0.021 0.023
GPT2-Large 774M 0.004 0.011 0.012 -0.004 -0.004 0.010 0.016
GPT2-XL 1.5B 0.036 0.016 -0.007 0.003 -0.0004 0.020 0.008
Effective Dimensionality
We computed the effective dimensionality of activation patterns using the participation ratio metric to assess functional specialization. Table 2 shows that rare token neuron groups consistently exhibit lower effective dimensionality compared to random baselines. This reduction in effective dimensions (boost: 33.0–43.0%, suppress: 32.2–43.0%) relative to random controls (36.2–46.0%) indicates more concentrated, specialized activation patterns within rare token neurons.
Table 2: Effective dimensionality proportions across neuron groups
Model Size Boost Suppress Random
Pythia-70M 70M 33.5 32.6 36.2
Pythia-410M 410M 33.0 32.2 37.3
GPT2-Small 124M 37.0 40.0 45.0
GPT2-Large 774M 43.0 43.0 46.0
GPT2-XL 1.5B 40.0 42.0 46.0
Cosine Similarity of Weight Vectors
Weight vector cosine similarities provide insight into the geometric organization of rare token neurons in parameter space. Table 3 reveals strong positive alignment within boost (0.028–0.141) and suppress (0.092–0.165) neuron groups, while showing consistent negative alignment between these specialized groups and random controls. The negative cross-correlations (boost vs. random: -0.100 to 0.003, suppress vs. random: -0.105 to -0.014) suggest that rare token neurons occupy distinct regions of weight space, supporting our hypothesis of functional specialization.
Table 3: Cosine similarity between weight vectors within and across neuron groups
Model Size Boost Suppress Random B vs. R S vs. R B vs. S R1 vs. R2
Pythia-70M 70M 0.141 0.165 0.021 -0.017 -0.014 0.146 0.021
Pythia-410M 410M 0.107 0.133 0.054 -0.032 -0.041 0.114 0.058
GPT2-Small 124M 0.109 0.122 0.089 -0.100 -0.105 0.120 0.099
GPT2-Large 774M 0.028 0.092 0.041 -0.034 -0.063 0.054 0.052
GPT2-XL 1.5B 0.095 0.095 0.009 -0.010 -0.015 0.090 0.012
Weight Distribution Analysis
We analyzed the heavy-tailed properties of weight distributions using power-law exponents ($\alpha$) estimated via the Hill estimator. Table 4 demonstrates that rare token neuron groups exhibit significantly lower $\alpha$ values compared to random controls, indicating heavier-tailed weight distributions. This finding supports our hypothesis that functional specialization correlates with the emergence of heavy-tailed statistical properties in neural networks.
Table 4: Power-law exponents ($\alpha$) for weight distributions across neuron groups
Model Size Boost Suppress Random
Pythia-70M 70M 4.30 3.97 6.37
Pythia-410M 410M 3.80 3.43 7.56
GPT2-Small 124M 2.12 1.57 6.74
GPT2-Large 774M 3.30 1.84 8.31
GPT2-XL 1.5B 2.01 1.68 9.33
Phase Transition Dynamics
Figure 4 illustrates the evolution of neuron influence distributions throughout training across the GPT-2 model family. The suppress neuron populations exhibit characteristic three-phase development: an initial plateau phase with uniform low influence, followed by a power-law scaling regime, and concluding with rapid decay at high influence values. This pattern emerges consistently across model scales, suggesting a universal mechanism underlying rare token neuron specialization.
Figure 4: Distribution of neuron influence slopes for suppress neurons across the GPT-2 model family, showing characteristic three-phase organization emerging during training.